Beta Company purchased machinery on June 1, 2018 for $89,000. It provided the following data:

Estimated Residual Value - $5 000

Estimated Service Life – 8 years

Estimated Total Work Hours – 42 000 hours

Estimated Total Production – 525 000 units

Machine Usage (2018) – 5 500 hrs

                             (2019) – 6 000 hrs

Machine Production (2018) – 48 000 units

                                      (2019) – 55 000 units

Beta Company uses a fractional-year policy to the nearest full month and its financial year ends December 31.

Required:

  1. What is the depreciable base of the machine? [1 mark]
  2. Calculate the depreciation expense for the year ending December 31, 2018 and 2019

Using the following methods: Round to nearest whole number

i. Straight-line. [3 marks]

ii. Units of output. [3 marks]

iii. Sum of the years’ digits. [5 marks]

iv. Double declining balance. [4 marks]
Answer #1

a ) Depreciable value = $ 89000 - $ 5000 = $ 84000.

i .Straight line Method

>> Depreciation per annum = ( $ 84000 / 8 ) = $ 10500.

>> Depreciation for 2018 = ($ 10500 * ( 7 / 12 ) )= $ 6125.

>> Depreciation for 2019 = $ 10500.

ii. Units of Output

Depreciation per unit = $ 84000 / 525000 = $ 0.16

>> Depreciation for 2018 = $ 0.16 * 48000 = $ 7680

>> Depreciation for 2019 = $ 0.16 * 55000 = $ 8800.

iii. Sum of years digit method

Sum of 8 years = 36.

Depreciation for 2018 = $ 84000 * ( 8 / 36 ) * 7 /12 = $ 10889.

Depreciation for 2019 = $ 84000 * ( 7.417 / 36 ) = $ 17306.

iv. Double Declining Method

Depreciation rate = ( 2 * 100 / 8 ) = 25 %

Depreciation for 2018 = $ 89000 * 25 % * ( 7 /12 ) = $ 12979.

Depreciation for 2019 = ( 89000 - 12979 ) * 25 % = $ 19005